UV-VISIBLE TRANSITION STRENGTHS OF THE HALOGENS IN THE GAS PHASE AND IN INERT LIQUID SOLVENTS

\underline{J. TELLINGHUISEN}, R. IAN GRAY, AND K. M. LUCKETT, Department of Chemistry, Vanderbilt University, Nashville, TN 37235.

The UV-visible absorption spectra of the homonuclear halogens involve three electronic bands – A-X, B-X, and B’(1_{u})-X. It has long been known that the absorption in inert solvents is stronger than that in the gas phase by 20-30I_{2} and Br_{2} in CCl_{4} and n-heptane are recorded at temperatures from 15 C to 50 C and are decomposed into component bands by least-squares analysis. The results of this analysis show: (1) the A-X band strength is the same within 5increased absorption strength in the liquid phase is attributable mostly if not entirely to the B’-X transition. The reasons for this behavior are under investigation.